OpenStudy (anonymous):

A parallel circuit has two 8.0-ohm resistors and a power source of 9.0 volts. If a 12.5-ohm resistor is added to the circuit in parallel, how will the current be affected and what value will it have? A. current will increase;2.97 B. current will increase;0.55 C. current will decrease;2.97 D. current will decrease;055

OpenStudy (agent0smith):

Find the total resistance first using this formula (since they're in parallel) \[\huge \frac{ 1 }{ R_T} = \frac{ 1 }{R_1 }+\frac{ 1 }{R_2 } + \frac{ 1 }{ R_3 }\] then just use V=IR to find the new current.

OpenStudy (anonymous):

so the answer is A

OpenStudy (agent0smith):

idk, i didn't work out the answer...

OpenStudy (agent0smith):

Current will increase tho, since adding resistors in parallel decreases total resistance.

OpenStudy (anonymous):

What is the current in a parallel circuit which has two resistors (17.2 ohms and 22.4 ohms) and a power source of 6.0 volts